root 3 sin theta - cos theta = 0 and 0 degree < theta < 90 degrees ?
Answer by stanbon(75887)

root 3 sin theta - cos theta = 0 and 0 degree < theta < 90 degrees ?
-----
3sin(t) - cos(t) = 0
----
Divide thru by cos(t) to get:
3tan(t) -1 = 0
tan(t) = 1/3
t = arctan(1/3) = 18.43 degrees
----------------------
Cheers,
Stan H.